OpenStudy (anonymous):

Tell me what I need to do next! Thank You! Given a cylinder with a surface area of 60 units squared and a height of 7 units. Find the volume of a sphere that has the same size radius as the given cylinder. Show all work. 2πr^2+2πrh = 60 pi 2pi(r^2 + rh)=60pi r^2+r•h=30 h =7 r^2+7•r=30 r(r+7) =30 if r = 3 then 3(3+7)=30 √ r=3

OpenStudy (asadkarim7):

the formula I use for surface area is universal fromula applies to all uniform shaped figures= perimeter of base x height x 2 ( base area)

OpenStudy (anonymous):

vol sphere (3/4)pi r^3

OpenStudy (asadkarim7):

2 pi r x 7 x 2 pi r ^2= 60 pi take pi common pi ( 2r x 7 x 2 r^2) = 60 pi eliminate pi 2r^2 x 2 r x 7 = 60 28 r ^3 = 60 r = cube root 2.14

OpenStudy (anonymous):

so i did it wrong?

OpenStudy (anonymous):

two first anwers tell main idea.

OpenStudy (anonymous):

vol sphere (4/3)*pi*R^3

OpenStudy (anonymous):

You showed your work, but I didn't really follow, but it appears you are finding the radius of cylinder. Once you have the radius, punch it in above sphere vol formula.

OpenStudy (anonymous):

I double-checked your work, 3 is correct radius. You want to find the volume of sphere of same radius\[(4/3)\pi3^{3}\]